### v1

# Federated Learning for Surgical Vision in Appendicitis Classification: Results of the FedSurg EndoVis 2024 Challenge

Abstract
Purpose: The FedSurg challenge was designed to benchmark the state of the art in federated learning for surgical video classification. Its goal was to assess how well current methods generalize to unseen clinical centers and adapt through local fine-tuning while enabling collaborative model development without sharing patient data.
Methods: Participants developed strategies to classify inflammation stages in appendicitis using a preliminary version of the multi-center Appendix300 video dataset. The challenge evaluated two tasks: generalization to an unseen center and center-specific adaptation after fine-tuning. Submitted approaches included foundation models with linear probing, metric learning with triplet loss, and various FL aggregation schemes (FedAvg, FedMedian, FedSAM). Performance was assessed using F1-score and Expected Cost, with ranking robustness evaluated via bootstrapping and statistical testing.
Results: In the generalization task, performance across centers was limited. In the adaptation task, all teams improved after fine-tuning, though ranking stability was low. The ViViT-based submission achieved the strongest overall performance. The challenge highlighted limitations in generalization, sensitivity to class imbalance, and difficulties in hyperparameter tuning in decentralized training, while spatiotemporal modeling and context-aware preprocessing emerged as promising strategies.
Conclusion: The FedSurg Challenge establishes the first benchmark for evaluating FL strategies in surgical video classification. Findings highlight the trade-off between local personalization and global robustness, and underscore the importance of architecture choice, preprocessing, and loss design. This benchmarking offers a reference point for future development of imbalance-aware, adaptive, and robust FL methods in clinical surgical AI.
keywords:
Federated Learning , EndoViS Challenge , Appendectomy , Video Classification , Surgical Data Science
journal: Medical Image Analysis
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
First federated learning challenge in surgical AI
First use of preliminary multi-center Appendix300 dataset
Novel patient-level video classification beyond image-based tasks
Findings on generalization vs. personalization in federated learning
Demonstrated strengths and limitations of federated surgical AI
1 Introduction
The combination of early successes in AI algorithms and the emerging field of Surgical Data Science (SDS) holds strong potential to transform surgery [1]. Recent work has demonstrated that AI can reliably analyze surgical video, comprehending anatomy, tool usage, and procedural events in real time [2]. Such capabilities are clinically relevant given the established relationship between video-derived quality indicators and postoperative complications. However, the development of robust AI systems critically depends on access to large volumes of high-quality, diverse surgical data. This requirement remains one of the field’s most pressing challenges [3].
Surgical datasets are typically sourced from either a single institution or a small group of collaborators [3, 4]. While these datasets enable preliminary developments, they inherently suffer from limited diversity and scale if applied in real-world scenarios [5]. Single-center datasets often lack the heterogeneity required for generalizable AI models and involve prolonged data acquisition cycles [6]. Multi-institutional data aggregation, on the other hand, can significantly enrich the dataset, but is frequently obstructed by stringent regulatory frameworks such as the Health Insurance Portability and Accountability Act (HIPAA) [7] or the General Data Protection Regulation (GDPR) [8] that prohibit unrestricted data sharing.
A viable solution is Federated Learning (FL). FL enables decentralized model training across multiple clients, such as hospitals, without the need to share raw image or video data [9]. Instead, the model is brought to the data, preserving privacy and complying with regulations like GDPR or HIPAA. In a standard FL workflow, each client trains a model locally on its private dataset. Only model updates are sent to a central server, where they are aggregated using strategies like Federated Averaging (FedAvg) [9] or more robust alternatives such as FedMedian [10]. The updated global model is then redistributed to clients, and this cycle repeats until convergence.
FL is especially well-suited for medical applications, where data privacy, security, and the lack of standardized datasets pose significant challenges [11]. It enables large-scale collaborative model development without the need to centralize sensitive, multi-institutional data, thereby preserving patient confidentiality and regulatory compliance [11]. However, FL also introduces new challenges. These include handling data and system heterogeneity across institutions [12, 13, 14], balancing personalization with generalization [15], managing communication overhead due to frequent model updates [9, 13], addressing fairness and potential model biases [13, 16], and coping with the increased complexity of evaluation and debugging in distributed settings [13]. In the context of SDS, recent studies have begun to explore FL for tasks such as surgical phase recognition, scene segmentation, and tool detection [17, 18, 19].
To address challenges in SDS, initiatives such as the Endoscopic Vision (EndoVis) Challenge [20] have become critical accelerators for progress. The Federated Learning for Surgical Vision (FedSurg) challenge, introduced as part of the 2024 edition of EndoVis, represents the first FL challenge in the field of SDS. It aimed to benchmark the state of the art in applying FL to surgical AI for privacy-preserving model development across institutions. In particular, the challenge focused on evaluating model performance in terms of generalization to unseen centers versus adaptation to individual centers, reflecting two core challenges of FL in real-world surgical applications. The Appendix300 dataset is a multi-institutional appendectomy video dataset [21]. As the first FL challenge of its kind in SDS, FedSurg pioneers the use of a preliminary version of the Appendix300 dataset as a foundational platform for algorithm development and validation. A separate benchmarking paper has been published for the complete dataset [22]. In this paper, we report the challenge design, results, and findings of FedSurg according to the transparent reporting of biomedical image analysis challenges (BIAS) guidelines [23].
The main contributions of this paper are:
  • 1.
    The design and results of FedSurg, the first international challenge in Surgical Data Science dedicated to Federated Learning, which pioneers the use of the preliminary, multi-center Appendix300 dataset.
  • 2.
    A benchmark of FL strategies for the novel task of patient-level surgical video appendicitis grading classification, advancing the complexity of analysis beyond common static, image-based tasks.
  • 3.
    A rigorous, data-driven analysis of the critical trade-off between model generalization to unseen institutions and personalization to local client data.
  • 4.
    A clear demonstration of the current strengths and practical limitations of FL in surgical AI, highlighted by bootstrapping and Wilcoxon signed-rank test.
2 Challenge Design
This section outlines the design of the FedSurg challenge, detailing its organizational structure, the core mission guiding the competition, and the datasets employed for evaluation. Additionally, it describes the assessment methods used to fairly and rigorously evaluate submitted models and FL settings, ensuring robust comparison across the multi-centric surgical video dataset.
2.1 Challenge Organization
The FedSurg challenge was a one-time event with a fixed submission deadline, held as part of the Medical Image Computing and Computer Assisted Intervention (MICCAI) 2024 conference in Marrakech, Morocco (October 6–10, 2024). Organized jointly by research groups from the Dresden University of Technology (TUD), Purdue University, and the National Center for Tumor Diseases (NCT) Dresden (see C), the challenge offered a prize pool of €500, provided by the Horizon Europe NearData project and distributed equally between the two challenge tasks.
The primary mission of the FedSurg challenge was to benchmark FL approaches for surgical video classification using a new multi-institutional dataset. The dataset was collected from four German hospitals under institutional ethical approval (see F). Only training data was released to participants, while test data remained private and was accessible only to the organizers. The dataset used in the FedSurg challenge is a preliminary subset of the Appendix300 dataset [21] (see Subsection 2.3).
All relevant information, including registration, data access, submission progress tracking, guidelines, and a discussion forum, was made available through the official challenge website on Synapse (see D). This platform served as the central hub for participant onboarding, communication, and submission management. Participants were required to sign challenge rules before they could participate and get access to the data (see B). The challenge timeline included registration in April 2024, a feedback-based testing phase during August–September, and the final submission deadline on September 18, 2024 (Figure 1).
Participants were restricted to the released training data and publicly available resources, including pre-trained models. Challenge-provided data could not be used for pre-training in FL submissions. Members of the organizing institutes were permitted to participate but were not eligible for awards. More information about the challenge is available in the challenge design document (see E).
Submissions were required in a containerized format (Docker) and were evaluated on a dedicated server equipped with up to 8 NVIDIA V100 GPUs, 56 CPUs, and 756 GB RAM. This standardized hardware environment and containerized execution ensured reproducibility and fair benchmarking under identical conditions (Figure 1). To support participants, example FL code and evaluation scripts were made publicly available (see D).
Results were first announced during the MICCAI 2024 Satellite Events and subsequently published on the challenge website, alongside the evaluation framework, rankings, and key performance metrics. All results and analyses from teams with a complete working submission are included in this joint publication, with contributing team members listed as co-authors. Authors were not permitted to publish individual challenge results prior to the release of this paper, ensuring coordinated dissemination and preserving the novelty of the outcomes.
Figure 1: FedSurg24 Challenge Highlights: The top panel shows example images of intraoperative appendicitis grades, defined according to Gomes et al. [24], which were used for video annotation. The lower panel illustrates the FedSurg Challenge workflow: teams submitted Docker containers via Synapse, which were executed on a secure cluster simulating FL across three centers with local training and centralized aggregation. Final performance was assessed by testing each center’s best local model on its own test set, while the global model was evaluated on the unseen hold-out center to measure generalization. The challenge timeline with key dates is shown alongside.
2.2 Challenge Mission
The FedSurg challenge focuses on classifying the inflammatory stage of acute appendicitis using laparoscopic appendectomy videos, an important and commonly performed surgical procedure. The primary objective is to provide a comparative benchmark of existing solutions, with a special emphasis on exploring various FL strategies. In particular, the challenge investigates the balance between personalization and generalization under inherent data heterogeneity. For this purpose, the newly created Appendix300 dataset was introduced, with the challenge of utilizing a partial subset of its data. This makes FedSurg not only the first FL challenge in SDS but also the first benchmarking study of different FL approaches on the patient-level task of appendicitis classification.
The challenge cohort comprises human patients, both adults and children of diverse ages and biological sex, who underwent appendectomy for suspected appendicitis at four German medical centers. In general, the challenge cohort is a representative sample of patients undergoing appendectomy for suspected appendicitis. The dataset consists of laparoscopic video recordings captured during these interventions. While no additional patient data was provided for the challenge, the final Appendix300 dataset includes supplementary clinical information [21].
This challenge’s main contribution is its patient-level prediction task for appendicitis staging, a transferable innovation for AI in surgery. By establishing a framework for standardized, objective assessment of inflammation severity, this approach serves as a powerful proof-of-concept for addressing more critical clinical needs. These include advancing intraoperative assistance for surgical quality control, and providing a foundation for developing AI tools for accurate diagnosis, real-time surgical support, and effective prognosis for a variety of surgical conditions.
Participants were tasked with developing FL algorithms to classify appendicitis stages in laparoscopic videos, structured around two core tasks:
  • 1.
    Task 1, Generalization: Evaluate the model’s ability to generalize to unseen centers. Participants train their models on a subset of centers and are evaluated on a held-out center that was not involved in training.
  • 2.
    Task 2, Adaptation: Assess the model’s ability to personalize to each center’s test data. Here, the same trained model is fine-tuned (or adapted) for each single center of the federated setup and then evaluated independently on each center’s test set.
In the context of federated learning, generalization is the ability of a collaborative global model to perform well on data from new clients, while personalization involves adapting the model to achieve optimal performance for a specific client’s unique data. The challenge promotes the creation of privacy-preserving algorithms that can tackle both of these issues, creating models that are both robust and adaptable to diverse surgical settings.
2.3 Challenge Dataset
The challenge dataset is a preliminary subset of the Appendix300 collection, comprising frames from 223 full-length recordings of laparoscopic appendectomies. Beyond image frames and laparoscopic grading, the finalized Appendix300 dataset is enriched with detailed histopathological findings and patient anamnesis data [21]. For this challenge, 200 frames were extracted per video using FFmpeg software [25], capturing a 100-second window sampled at two frames per second around an annotated timestamp [21].
Frames were selected at the timestamp identified by the operating surgeon when the appendix was fully visible prior to dissection [21]. Inflammation severity, graded from level 0 to 5, was annotated by the operating surgeon (surgery residents with varying years of experience). The annotation protocol (available at [21]) defines class descriptions and includes illustrative examples based on the definition by Gomes et al. [24]. Furthermore, fine-grained classes 3A/3B were merged into class 3, and classes 4A/4B into class 4 in this challenge dataset. An example overview of the data is visible in Figure 1. By using this annotation protocol, a verbal explanation to the participating surgeons, and a custom graphical user interface of the annotation software, we minimize misclassification. Therefore, we deem the risk of misclassification limited to borderline cases (i.e., cases between partial and total necrosis of the appendix). This potential source of error applies equally to all centers.
The challenge dataset comprises data collected from multiple hospitals across Germany, including both university and community hospitals with varied surgical profiles. While the challenge simulates an FL setup, it reflects a realistic scenario since the data originates from four distinct, real-world centers that have been anonymized to protect institutional privacy. The contributing institutions include:
  • 1.
    Asklepios-ASB Klinik Radeberg
  • 2.
    University Hospital Carl Gustav Carus Dresden, Department of Pediatric Surgery
  • 3.
    Krankenhaus St. Joseph-Stift, Dresden
  • 4.
    St. Elisabethen-Krankenhaus, Ravensburg
An overview of the video distribution across centers is shown in Table 1.
Table 1: Data Distribution Across Participating Centers.
Center Training Videos Testing Videos Total
1 40 10 50
2 33 9 42
3 80 22 102
4 0 29 29
Total 153 70 223
(a) Training Dataset Distribution.
(b) Testing Dataset Distribution.
Figure 2: Label Distribution Across Data Subsets per Center. Label distributions for (a) the training dataset and (b) the test dataset across the four centers. The plots highlight notable inter-center variability and class imbalance. In the training set visualization, the darker segments represent the publicly available subset for participant development, while the lighter segments show the complete dataset used for final federated training. The exact data distribution was unknown for the participants.
To discourage centralized model development and preserve the integrity of the federated setting, the training data was partitioned into public and private subsets. Participants were granted access only to the public subset, which constituted 25% of the total training data. The remaining 75% was used exclusively by the organizers during the final training phase, where both public and private data were jointly leveraged in a secure federated setup (Figure 1 and Figure 2). The class distribution within the dataset mirrors real-world clinical prevalence, featuring a predominance of mid-level cases and a lower representation of extreme or mild inflammation levels.
No predefined validation set was provided. Participants were expected to devise their own validation strategies. Final submissions were executed centrally by the organizers, who conducted full federated training and evaluation across all centers under secure and standardized conditions to ensure reproducibility and fairness.
Since the challenge dataset constitutes a preliminary subset of the Appendix300 dataset, we provide a CSV file specifying which samples were used in the challenge here [^1] Available upon acceptance or on request.. In addition, two videos are released that were excluded from the final dataset because they were either shorter than 100 seconds or did not contain a clearly visible appendix in the center. The Appendix300 dataset is publicly available for non-commercial use under the Creative Commons Attribution (CC BY) license. Any use of this dataset requires citing both this paper and the final Appendix300 publication.
2.4 Assessment Methods
The evaluation of this challenge is based on two complementary metrics: the macro-averaged F1-score, also known as the Dice coefficient, and the Expected Cost (EC) with linear weights. Together, these metrics provide a robust assessment of model performance in a multi-class setting with ordinal structure.
F1-Score (Dice Coefficient):
The F1-score is widely used for balancing precision and recall, offering a harmonic mean that reflects both the accurate identification of relevant instances and the minimization of false positives [26, 27]. Given a confusion matrix $\mathbf{M}\in\mathbb{N}^{C\times C}$, where $M_{i,j}$ denotes the number of samples with ground-truth class $i$ predicted as class $j$, the F1-score for class $c$ is computed as:
$\text{F1}_{c}=\frac{2\cdot\text{TP}_{c}}{2\cdot\text{TP}_{c}+\text{FP}_{c}+\text{FN}_{c}}$ (1)
where:
$\text{TP}_{c}$ $=M_{c,c}\quad\text{(true positives)}$
$\text{FP}_{c}$ $=\sum_{\begin{subarray}{c}i=1\\ i\neq c\end{subarray}}^{C}M_{i,c}\quad\text{(false positives)}$
$\text{FN}_{c}$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq c\end{subarray}}^{C}M_{c,j}\quad\text{(false negatives)}$
The overall F1-score is calculated as the macro-average across all classes:
$\text{F1}_{\text{macro}}=\frac{1}{C}\sum_{c=1}^{C}\text{F1}_{c}$ (2)
Expected Cost (EC)
To account for the ordinal nature of the classification task, we also report the Expected Cost (EC), which penalizes misclassifications based on their severity. This aligns with the principle of ordinal monotonicity, where predictions farther from the ground-truth class incur higher penalties [28, 29, 27]. The EC is defined as:
$\text{EC}=\frac{1}{N}\sum_{i=1}^{C}\sum_{j=1}^{C}M_{i,j}\cdot w_{i,j}$ (3)
where:
$N$ $=\sum_{i=1}^{C}\sum_{j=1}^{C}M_{i,j}$ (total number of samples)
$w_{i,j}$ $=\frac{|i-j|}{C-1}$ (cost of predicting class $j$ when the true class is $i$)
with the linear weight function:
$w_{i,j}=\frac{|i-j|}{C-1}$ (4)
This assigns zero cost to correct predictions and a linearly increasing penalty for deviations, with a maximum cost of 1 for the farthest misclassifications.
2.4.1 Ranking
To ensure a fair evaluation of submissions across both generalization and adaptation scenarios, we implemented a ranking framework that integrates multiple performance metrics with rigorous statistical analysis. Only submissions that successfully completed both tasks were considered. Teams failing to meet this criterion or submitting non-executable code were disqualified. Our ranking methodology guaranteed that models were evaluated not just on average performance but also on their robustness and consistency across heterogeneous data settings. The details of our evaluation procedure are described below.
For Task 1 (generalization), we computed the F1-score and the EC metric for all test cases. Separate rankings were assigned based on each metric, and a team’s overall rank for Task 1 was derived by averaging these two ranks. For Task 2 (adaptation), the metrics were first averaged across all three centers, and the same ranking scheme used in Task 1 was then applied.
The final leaderboard was determined by averaging each team’s ranks across both tasks, providing a comprehensive assessment of overall performance throughout the challenge.
Additionally, to ensure the robustness and reliability of the rankings, bootstrapping [30] was applied. Bootstrapping, as emphasized by Maier-Hein et al. [31], is a key method for assessing the variability of rankings and the stability of observed performance differences. In this study, the test set was repeatedly resampled with replacement for 10,000 iterations, and team rankings were recalculated for each resample. For each team, we computed the proportion of iterations in which it retained its original rank as well as the proportions in which it achieved each of the other possible ranks. To statistically compare team performances, we applied the Wilcoxon signed-rank test to the bootstrapped metric values obtained over all iterations, thereby quantifying whether observed performance differences were significant beyond random variation.
3 Results
The FedSurg challenge received 24 registrations, with four final submissions. However, one submission was disqualified due to non-executable code after the final deadline. This left three complete submissions for evaluation. The participating teams were: Santhi R. Kolamuri, who submitted independently as Team Santhi; Lorenzo Mazana and Claas de Boer from the Translational Surgical Oncology group at the National Center for Tumor Diseases, submitting as Team Elbflorenz; and Julia Alekseenko and Nicolas Padoy from the CAMMA research group at IHU Strasbourg, submitting as Team Camma.
3.1 Participating Teams and Methods
The following section details the methodologies submitted by participants in the FedSurg challenge. Each subsection outlines the architectural choices, training strategies, and FL configurations employed. Where relevant, we contextualize each approach within existing literature on foundation models, metric learning, and federated optimization. A summary of each submission is presented in Table 2.
Figure 3: Methods Overview: The three submissions shown utilize different backbone architectures and federated strategies. A common approach is that in each server round, the best-performing model from a client’s local training rounds is sent to the server for aggregation. (a) Team Santhi uses a frozen ViViT backbone with a fine-tuned classification head processing 32 frames per video, with updates aggregated via FedAvg. (b) Team Elbflorenz uses a frozen EndoViT backbone with a fine-tuned head, predicting single frames repeatedly and combining them via majority voting, with updates aggregated via FedSAM. (c) Team Camma uses ResNet50 models trained with a contrastive approach on positive and negative pairs, with updates aggregated via FedMedian. At inference, classification is performed by comparing the test embedding to a support set.
Table 2: Overview of submissions to the FedSurg Challenge, detailing each team’s backbone model, prediction method, frame sampling, loss function, optimizer, learning rate, batch size, and federated learning configuration (FL strategy, FL-rounds, and local rounds).
Team Santhi Elbflorenz Camma
Backbone Model Pretrained Video Vision Transformer (ViViT) with a frozen backbone. Pretrained EndoViT with a frozen encoder. Siamese network with a ResNet-50 backbone.
Prediction Method Video-level classification. Majority voting on independent frame-level classifications. Compares test embedding to class prototypes or a full support set.
Frame Sampling Samples 32 frames, with two-thirds from a narrow window around the center and the rest from the full sequence with a bias towards the center. Uses 100 equidistant frames from each video. Selects representative frames based on cosine similarity to a reference embedding, plus the central keyframe.
Loss Function Cross-entropy loss. Weighted cross-entropy loss to address class imbalance. Triplet margin loss with cosine similarity.
FL Strategy Federated Averaging (FedAvg). Adaptive Federated Sharpness-Aware Minimization (FedSAM) on Client-side and FedOpt on Server Side. Federated Median (FedMedian).
Optimizer Adam adaptive Federated Sharpness Aware Minimization (FedSAM) based on SGD Adam
FL-Rounds 5 50 10
Local Rounds 20 2 5
Learning Rate $\times 10^{-4} 1×1$$\times 10^{-3}$ $1\times 10^{-6}$
Batchsize 4 128 1
3.1.1 Team Santhi
Santhi R. Kolamuri’s approach leverages a pretrained Video Vision Transformer (ViViT) [32], which is well-suited for capturing spatio-temporal features in surgical video sequences (Figure 3 (a)). ViViT’s transformer-based architecture has proven more effective than conventional CNNs in modeling temporal dependencies. By utilizing ViViT pretrained on Kinetics-400, only the final classification layer is fine-tuned, while all other weights remain frozen. This linear probing approach is common for foundation models [33], as it allows efficient adaptation while retaining robust pretrained representations.
The frame loading strategy is customized to emphasize salient temporal regions. From 200 available video frames, 32 are sampled for training. Two-thirds are selected from a narrow window around frame 100 to capture high-information segments, while the remainder of the frames are drawn from the full sequence with 60% probability to the frames near the center. This hybrid sampling balances local relevance and temporal diversity.
Training is implemented via the Flower FL framework [34], using FedAvg for aggregation. Centers train locally with a batch size of four for five epochs per round, across 20 communication rounds. Cross-entropy loss is used with mixed precision training to improve memory efficiency. The lightweight design, frozen backbone, focused sampling, and shallow head enable fast convergence with low resource demands. This submission builds on the success of ViViT in centralized video classification while adapting it to the constraints of FL.
3.1.2 Team Elbflorenz
Team Elbflorenz adopts a foundation model-based strategy using EndoViT [35], a Vision Transformer pretrained on the Endo700k dataset consisting of diverse endoscopic images (Figure 3 (b)). Following a standard linear probing setup [33], the encoder is frozen and only a lightweight classification head is trained. This setup is motivated by recent successes of foundation models in medical imaging [36, 37], where pretrained encoders generalize well even with limited task-specific data.
The model uses 100 equidistant frames from each video and independently classifies each frame. A weighted cross-entropy loss is used to mitigate class imbalance, with weights inversely proportional to class frequency. Final video-level predictions are determined via majority voting. In the event of a tie, average confidence scores guide the decision. This per-frame classification strategy increases robustness by aggregating multiple frame-level predictions.
Federated optimization on the client side is handled using adaptive FedSAM [38, 39], which extends Sharpness-Aware Minimization (SAM) to heterogeneous FL settings. This method encourages flatter minima and better generalization across non-IID (independent and identically distributed) client data. Similar to Team Santhi’s approach, training was implemented via the Flower FL framework [34]. It employs 2 local epochs and 50 global rounds, with center-side class balancing and hyperparameter tuning. Compared to standard FedAvg, FedSAM has demonstrated improved convergence on non-IID data [38]. Model aggregation on the server is done with FedOpt [40]. By combining EndoViT with adaptive optimization, Team Elbflorenz presents a minimal yet effective pipeline for surgical appendicitis classification under federated constraints.
3.1.3 Team Camma
Team Camma employs a metric learning approach based on a Siamese network with a ResNet-50 backbone [41], inspired by the original Siamese architecture [42]. Unlike conventional classifiers that output class probabilities, this model maps input images to L2-normalized 256-dimensional embeddings. Training is performed using triplet margin loss with cosine similarity, encouraging embeddings from the same appendicitis stage to cluster while pushing apart those from different stages (Figure 3 (c)).
At inference, classification is performed by comparing the test embedding to a support set. This is done either via class prototypes (mean embeddings) or by averaging distances to all embeddings within each class. Camma observed that prototype-based inference worked best for Centers 2 and 3, while per-sample comparison was more effective for Center 1, allowing the approach to flexibly adapt to center-specific data distributions.
This metric learning paradigm is particularly well-suited for FL, as it focuses on learning a robust embedding space rather than simply aggregating classifier weights, which can be highly sensitive to the non-IID label distributions across centers.
To address domain heterogeneity, the model incorporates Switchable Normalization [43], which combines batch, instance, and layer normalization through softmax gating. This adaptive scheme enhances generalization across diverse centers, which is crucial in federated settings.
In addition, frame selection leverages embedding-based similarity: representative frames are chosen based on cosine similarity to a ResNet-50 reference embedding, alongside the keyframe (frame 100), to improve input consistency.
Training is conducted within a custom FL framework using the FedMedian aggregation strategy [10], which is robust to outliers. Each center performs 10 local epochs followed by 5 epochs dedicated to triplet optimization. The locally best-performing model (by F1-score) is selected for aggregation, ensuring only high-quality updates contribute to the global model. Overall, Team Camma’s method demonstrates how metric learning, adaptive normalization, and robust aggregation can be combined for scalable, personalized FL in surgical video classification.
3.2 Scores and Rankings
The performance of the submitted federated models was evaluated in both tasks of the FedSurg Challenge: generalization to an unseen center (Task 1) and center-specific adaptation (Task 2). Performance was measured using Expected Cost (EC) and F1-score. Rank stability was assessed via 10,000-iteration bootstrapping, followed by the Wilcoxon signed-rank test for statistical significance. A comprehensive summary of results—including key metrics, confusion matrices, and rank stability—is provided in Tables 3 and 4 and Figures 4, 5, 6, and 7, with further details in Appendix G.
3.2.1 Task 1: Generalization to an Unseen Center
Table 3: Performance comparison of the three teams on Task 1 at the held-out center (Center 4), reporting Expected Cost (EC, lower is better) and F1-Score (higher is better). Best results are highlighted in bold.
Team EC $\downarrow$ F1 $\uparrow$
Camma 57.24% 4.76%
Elbflorenz 24.14% 7.83%
Santhi 12.41% 23.03%
Figure 4: Confusion Matrices – Task 1, Center 4. Confusion matrices for the participating teams on Center 4 (Task 1). The values in the confusion matrices are not normalized. The color highlighting is normalized row-wise by true labels. The diagonal highlights class-wise recall, while off-diagonal values indicate common misclassification patterns.
Figure 5: Bootstrapped Performance Results. Visualization of the performance results with standard deviation as error bars for all teams and tasks after bootstrapping with 10,000 repetitions. The plot illustrates the variability and stability of the outcomes across different centers.
Figure 6: Ranking Stability. Bootstrapped ranking distributions for each metric and task, based on 10,000 bootstrap iterations. Circle size indicates the percentage of times a team’s model achieved a specific rank across samples. Black crosses show median ranks, and black lines denote the 95% bootstrap confidence intervals. Subfigures (a) and (b) correspond to Task 1 (generalization ability) with metrics EC and F1-score, respectively, while (c) and (d) represent Task 2 (adaptation ability) using the same metrics.
Task 1 assessed the models’ ability to generalize to Center 4, which was excluded during training. Overall, performance on this task was modest across all teams (Table 3).
Team Santhi achieved the highest F1-score (23.03%) and the lowest EC (12.41%), demonstrating superior generalization under domain shift. This is evident in the confusion matrix (Figure 4), where Santhi’s predictions tend to align with the diagonal, in contrast to the predictions heavily biased towards a specific class. The model correctly identified class 3 in 8 out of 11 cases but struggled with adjacent stages (e.g., classes 2 and 4) and underrepresented ones (e.g., 0, 1, 5), likely due to class imbalance. Bootstrapped performance metrics (Figure 5) confirm the statistical robustness of these findings. This is further supported by the ranking stability plots (Figures 6(a) and 6(b)), where Team Santhi ranks first in 99.23% of bootstrapped samples for F1-score and 99.79% for EC.
Team Camma obtained the lowest F1-score (4.76%) and the highest EC (57.24%). The model predominantly predicted class 0, achieving only three correct predictions across the dataset. This behavior indicates a failure to generalize, likely caused by the model overfitting to the training distribution and being unable to adapt to the significant domain shift in the unseen test set. The poor performance of this trivial classifier is supported by bootstrapping analysis (Figure 5) and statistical significance testing (Table 3). Across 10,000 repetitions, Team Camma consistently ranked as the lowest-performing team in 100% of cases for the EC-Score and 75.09% for the F1-score (Figure 6).
Team Elbflorenz also demonstrated limited performance, predominantly predicting class 2 across inputs (Figure 4). The model achieved an EC of 24.14% and an F1-score of 7.83%, ranking slightly above Team Camma. This is likely due to the classifier exhibiting a strong bias towards predicting classes adjacent to the most frequently observed labels in the training data. Confusion matrices and bootstrapped scores confirm weak generalization, placing Elbflorenz in the middle of the three submissions in terms of overall performance (Figure 6, Figure 5).
Wilcoxon signed-rank tests confirmed statistically significant performance differences between all models at Center 4. This validates that the observed variations in EC and F1-score reflect meaningful distinctions in generalization ability under domain shift. The ranking stability plot (Figure 6) visually reinforces this trend, clearly showing Team Santhi in the lead, followed by Elbflorenz and Camma. Therefore, the leaderboard presented in Table 5 is robust.
3.2.2 Task 2: Center-Specific Adaptation
Table 4: Performance comparison of the three teams on Task 2 across Centers 1, 2, and 3. The table reports Expected Cost (EC, lower is better) and F1-Score (higher is better). Best results per metric and center are highlighted in bold.
Team Center EC $\downarrow$ F1 $\uparrow$
Camma 26.67% 3.70%
Elbflorenz 1 17.78% 20.20%
Santhi 22.22% 20.83%
Camma 17.78% 30.28%
Elbflorenz 2 24.44% 3.70%
Santhi 20.00% 13.33%
Camma 20.91% 22.76%
Elbflorenz 3 21.82% 15.51%
Santhi 19.09% 12.04%
Camma 21.79% 18.91%
Elbflorenz Average 21.35% 13.14%
Santhi 20.44% 15.40%
Figure 7: Confusion Matrices – Task 2, Centers 1–3. The values in the confusion matrices are not normalized. The color highlighting is normalized row-wise by true labels. The diagonal highlights class-wise recall, while off-diagonal values indicate common misclassification patterns.
Task 2 evaluated model performance in intra-center settings, where each local model was tested on its dedicated test set from its associated center. All teams demonstrated improvements over Task 1, although performance varied considerably across centers (Table 4, Figure 5).
Team Santhi demonstrated robust performance across centers, with F1-scores of 20.83% (Center 1), 13.33% (Center 2), and 12.04% (Center 3) (Table 4). Although not consistently the best at individual centers, their steady results, combined with relatively low EC values, indicate effective regularization and strong adaptability to the test data of the training centers. The confusion matrices in Figure 7 reveal predictions largely concentrated along the diagonal, suggesting good class differentiation. In Center 3, however, the local model—similar to Team Elbflorenz—showed limited discriminatory power, with a tendency to predict classes 2, 3, or 4 rather than providing balanced classification. Bootstrapping analysis (Figure 5) confirms consistent performance across centers, but the final ranking remained unstable, as indicated by the wide 95% confidence intervals in Figure 6. For the F1-score, Santhi ranked first in 17.86% of iterations, second in 45.02%, and third in 37.12%; for EC, the respective values were 44.12%, 33.94%, and 21.94%.
Team Elbflorenz displayed greater inconsistency across centers, especially in the F1-Score (Figure 5). The confusion matrices (Figure 7) reveal inconsistent local adaptation: the models for Centers 1 and 2 collapsed into trivial classifiers, biasing predictions towards a single class, while the model for Center 3 failed to discriminate effectively between classes 2, 3, and 4. Despite this, the model performed relatively well at Center 1 (F1: 20.20%), but performance dropped markedly at Center 2 (3.70%) and moderately at Center 3 (15.51%). Bootstrapping results suggest difficulties in achieving reliable within-center generalization, potentially attributable to data imbalance or overfitting. Similar to Team Santhi, Team Elbflorenz’s ranking fluctuates between first, second, and third place across evaluation metrics (Figure 6). For the F1-Score, Elbflorenz ranked first in 10.23% of cases, second in 35.03%, and third in 54.74%. Regarding the EC metric, Elbflorenz achieved first place in 44.12% of cases, second place in 33.94%, and third place in 21.94%.
Team Camma exhibited pronounced center-dependent performance (Table 4, Figure 5), excelling at Center 2 with an F1-score of 30.28% and at Center 3 with 22.76%, but underperforming significantly at Center 1, where the F1-score was 3.70%. Confusion matrices shown in Figure 7 indicate intermediate performance relative to Teams Santhi and Elbflorenz, as the models demonstrate a modest ability to differentiate between classes, evidenced by a visible diagonal tendency. Similar to the other submissions, bootstrapping analyses confirm variability in performance across centers. The ranking for Team Camma also fluctuates across centers, with F1-score rankings of 1st place in 71.29% of cases, 2nd place in 19.94%, and 3rd place in 8.14%. For the EC metric, the rankings were 1st place in 27.81%, 2nd place in 29.55%, and 3rd place in 42.64%.
Importantly, Wilcoxon signed-rank tests revealed statistically significant performance differences between the models for all centers and both metrics. Conducting the test after bootstrapping substantially supports the statistical power, ensuring that the detected differences were not due to chance. Despite these significant differences, the bootstrapped ranking stability analysis (Figures 6(c) and 6(d)) shows notable fluctuations in team positions within the bootstrapping. While the Task 2 F1-score rankings (Figure 6(d)) align with the overall standings in Table 5, the EC mean-based rankings (Figure 6(c)) reveal no clear winner with respect to the median rank. Nevertheless, the frequency of rank positions matches the trend of the general ranking in Table 5. Such variability highlights that Task 2 outcomes are sensitive to small changes in the test set, and claims regarding top-performing submissions should therefore be made with caution. In line with the interpretation of Maier-Hein et al. [31], the fact that no team retained its original EC rank in at least 50% of bootstrap replicates indicates that the Task 2 ranking is unstable. We therefore report the original ranking for completeness, while emphasizing that it should not be interpreted as conclusive evidence of submission superiority.
Table 5: Team rankings for the FedSurg Challenge are presented, where lower ranks indicate better performance.
Team Task 1 Ranks Task 2 Ranks Final Rank
EC F1-Score Avg EC F1-Score Avg
Camma 3 3 3 3 1 2 2
Elbflorenz 2 2 2 2 3 3 2
Santhi 1 1 1 1 2 1 1
4 Discussion
4.1 Task Wise performance
The FedSurg Challenge establishes the first benchmark for evaluating Federated Learning (FL) strategies in the surgical video classification of appendicitis grades, utilizing a heterogeneous, multi-institutional dataset. Three complete submissions were received and evaluated on a preliminary subset of the Appendix300 dataset.
Task 1, which focused on generalization to an unseen center, immediately revealed the limited generalization abilities of current FL models. This difficulty stems directly from significant data heterogeneity across the participating institutions. For instance, images originated from a wide array of surgical video systems, including Arthrex Synergy UHD4, Storz image1 s, B. Braun Aesculap EinsteinVision, and Richard Wolf Endocam 4k, leading to variations in lighting and resolution [21]. For example, Center 1 provided 4k images, while Center 4 provided cropped images in Full HD resolution. In addition to this feature skew, a label distribution skew existed between the different centers (Figure 2). In this challenging environment, models capable of temporal modeling, like the ViViT-based submission, achieved the strongest but still weak results. A slight tendency also emerged showing that approaches focused on the loss landscape, like FedSAM, performed slightly better than a FedMedian aggregation of a triplet margin loss and cosine similarity model. This aligns with the conclusions of Foret et al. [39] and Caldarola et al. [38] that the quest for flat minima, which is incorporated into FedSAM, can lead to better generalization ability. This difficulty was not unique to the challenge submissions; a benchmark study on the Appendix300 dataset [22] using a leave-one-out cross-validation design similarly confirmed the profound challenge of inter-center generalization.
Task 2, which involved center-specific adaptation after fine-tuning, led to only marginally better results across all teams. This marginal improvement, however, came at the cost of rank instability, highlighting that even minor gains in local adaptation can compromise the robustness of a model’s comparative performance. The limited improvement can be attributed to several factors inherent to the federated setup. Due to the difficult knowledge of the complete data distribution and amounts, hyperparameter tuning was challenging, which could lead to local models overfitting to their local training data and, therefore, still failing to achieve high accuracy. Furthermore, the knowledge from other centers within the global model can negatively influence client performance. At each training round, the client initializes its weights from the global model, which represents an aggregated compromise of the knowledge learned across all participating centers. The poor local performance could stem from conflicting updates that pull the model away from what is optimal for its own center’s data.
Ultimately, these findings illustrate the central challenge of balancing generalization and adaptation, alongside the overarching difficulty of data heterogeneity in federated setups.
4.2 Stability assessment
The bootstrapped rank frequency analysis yielded distinct outcomes for Task 1 and Task 2.
For Task 1, the analysis indicated stable leader board rankings, with the predictions from submitted models being statistically distinct. A central finding was the critical role of temporal information; the highest-ranking submission utilized a video-based model, whereas models predicated on static frames exhibited the poorest performance. This aligns with the findings from the Appendix300 benchmarking paper [22], where the model trained only on middle frames performed poorly in comparison to the models trained on video models with a temporal horizon. Furthermore, using a higher frame rate up to 1 frame per second (fps) resulted into better results.
In stark contrast, the rankings for Task 2 demonstrated pronounced instability. No submission maintained its rank on the EC metric in more than 50% of the bootstrap replicates, indicating that the leader board outcomes were not robust. This instability suggests that leader boards for federated adaptation tasks should be interpreted with extreme caution, as top-ranking methods may not be reliably superior but merely fortunate in a specific data configuration.
4.3 Systemic Limitations and Data-Related Challenges in Federated Surgical AI
Several systemic issues emerged. FL does not resolve underlying dataset problems such as class imbalance and can amplify heterogeneity effects. Rare inflammation stages were underrepresented, leading to poor classification across all teams. Personalized losses and fine-tuning improved performance within centers (e.g., Team Camma) but compromised generalization to unseen centers, underscoring a fundamental trade-off in FL. This mirrors our findings from the primary tasks, where the unstable gains in center-specific adaptation (Task 2) failed to translate into the robust, generalizable models needed for unseen centers (Task 1). Similarly, design choices impacted outcomes: lightweight frozen foundation models provided stability across centers, whereas more adaptive approaches (e.g., Siamese networks with triplet loss) improved local performance but struggled with generalization under distribution shifts.
Furthermore, hyperparameter tuning proved difficult without access to the true data distribution, reflecting real-world FL constraints. The unknown class distribution contributed to lower-than-expected performance. Moreover, human experts showed only moderate agreement on inflammation grading ($\kappa_{weighted}=0.62$ [21]), underscoring the inherent ambiguity of the classification task.
4.4 Future directions
Future work should address identified limitations. Imbalance-aware strategies, such as reweighting, resampling, or federated focal loss, are needed to improve performance on clinically critical minority classes. Context-aware preprocessing and targeted frame sampling consistently outperformed uniform strategies and should be further explored. Methodological advances such as federated domain adaptation, personalized aggregation, and uncertainty-aware inference may improve robustness and safety. Finally, self-supervised pretraining on large-scale surgical video datasets could yield more transferable representations without compromising data privacy.
4.5 Challenge insights
Despite 24 registered teams, only three complete submissions were ultimately received. This discrepancy reflects both the niche character of FL within surgical AI and the complexity of the challenge setup. Unlike conventional challenges, FedSurg required fully encapsulated end-to-end submissions in Docker, with limited access to data and restricted debugging capabilities. These hurdles highlight practical barriers to broader participation and point to the technical demands of real-world FL applications.
5 Conclusion
The FedSurg Challenge provides the first benchmark for FL in surgical video classification, highlighting both opportunities and current limitations. The findings demonstrate that while spatiotemporal modeling and preprocessing choices can support generalization and adaptation, challenges persist in handling data imbalance, hyperparameter tuning, and stability of results. By exposing these trade-offs, FedSurg establishes an important reference point for the surgical AI community and guides future work toward robust, adaptive, and clinically relevant FL methods.
Acknowledgments
This work was co-funded by the European Union through NEARDATA under grant agreement ID 101092644, the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Dresden University of Technology, and the Federal Ministry of Education and Research of Germany in the program of “Souverän. Digital. Vernetzt.”, a joint project 6G-life with the project identification number 16KISK001K.
Furthermore, FRK receives support from the German Cancer Research Center (CoBot 2.0), the Joachim Herz Foundation (Add-On Fellowship for Interdisciplinary Life Science), the Central Indiana Corporate Partnership AnalytiXIN Initiative, the Evan and Sue Ann Werling Pancreatic Cancer Research Fund, and the Indiana Clinical and Translational Sciences Institute (EPAR4157) funded, in part, by Grant Number UM1TR004402 from the National Institutes of Health, National Center for Advancing Translational Sciences, Clinical and Translational Sciences Award. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
Team CAMMA (NP, JA) was supported by French State Funds managed by the Agence Nationale de la Recherche (ANR) under Grants ANR-22-FAI1-0001 (Project DAIOR) and ANR-10-IAHU-02 (IHU Strasbourg).
Conflict of Interest
JNK declares consulting services for Bioptimus, France; Panakeia, UK; AstraZeneca, UK; and MultiplexDx, Slovakia. Furthermore, he holds shares in StratifAI, Germany, Synagen, Germany, Ignition Lab, Germany; has received an institutional research grant by GSK; and has received honoraria by AstraZeneca, Bayer, Daiichi Sankyo, Eisai, Janssen, Merck, MSD, BMS, Roche, Pfizer, and Fresenius.
FRK declares advisory roles for Radical Healthcare, USA; and the Surgical Data Science Collective, USA, and has received research funding from Novartis.
SS received speaker’s fees from Stryker Corporation.
AR, NP, SS, DS, and SBa serve as an Associate Editor for the Medical Image Analysis journal.
NP is co-founder and owns shares in Scialytics SAS.
Declaration of generative AI and AI-assisted technologies in the writing process.
During the preparation of this work, the author(s) used ChatGPT, Gemini 2.5 Pro and DeepL Write in order to improve the readability and language of the manuscript. After using these tools/services, the author(s) reviewed and edited the content as needed and take full responsibility for the content of the published article.
Appendix A Dataset and Annotation Protocol
The challenge dataset is a preliminary subset of the publicly available Appendix300 dataset [21]. A CSV file detailing the samples used in the challenge is available here  [^2] Available upon acceptance or on request., along with two videos that were excluded from the final dataset but used in the challenge for reproducibility. The dataset is available for non-commercial use under a CC BY license and requires citation of the Appendix300 publication. The annotation protocol is available in [21]. Usage of the data for individual publications was prohibited before the release of this study and the Appendix300 dataset.
Appendix B Challenge Rules
During registration, participants signed the EndoVis rules document [^3] Available at https://tinyurl.com/nhb5z6a3.
Appendix C Challenge Organization
This work was jointly organized by four parties. Fiona R. Kolbinger, from the Department of Visceral, Thoracic and Vascular Surgery, Faculty of Medicine and University Hospital Carl Gustav Carus of the TUD Dresden University of Technology and the Weldon School of Biomedical Engineering, Purdue University coordinated the challenge data collection from four hospitals in Germany. Oliver S. Lestner and Jakob N. Kather, also from the Else Kröner Fresenius Center for Digital Health, supported the data aggregation from the technical point of view and obtained the source data from these centers. Max Kirchner, Alexander C. Jenke, Sebastian Bodenstedt, and Stefanie Speidel, from the Department of Translational Surgical Oncology (TSO) of the National Center for Tumor Diseases (NCT) Dresden, supported data aggregation and carried out the challenge organization and the technical implementation, including preprocessing, data provision, participant registration and administration, submission handling, evaluation, and results presentation. The NearData Horizon Europe program provided €500 in prize money, equally distributed across both tasks. If a single team achieved the top score in both tasks, they were eligible to receive both awards.
Appendix D Submission Instructions
The submission process was described in detail on the official challenge website [^4] Available at https://www.synapse.org/Synapse:syn53137385/wiki/625370. To support development, an example setup based on the FL Flower framework [34] was provided through a GitLab repository [^5] Available at https://gitlab.com/nct_tso_public/challenges/miccai2024/FedSurg24. The evaluation framework was made transparent through another repository [^6] Available at https://gitlab.com/nct_tso_public/challenges/miccai2024/snippet, which included the source code for metric computation as well as the ranking scripts (bootstrapping, statistical testing, and related plots).
Participants submitted their solutions as Docker containers via the challenge website. Docker Compose scripts with additional code were also accepted, provided they encapsulated the complete FL algorithm for both training and inference. Submissions had to run fully automatically without user interaction. Participants received email notifications about submission status and were allowed unlimited resubmissions until the final deadline. While the organizers do not distribute Docker images, teams were encouraged to release their code publicly.
Appendix E Challenge design document
See Supplementary file S1.
Appendix F Ethics approval
This study was prospectively reviewed and approved by the Institutional Review Board of the TUD Dresden University of Technology, Germany (approval number: BO-EK-332072022, approval date: August 4, 2022). The corresponding study was prospectively registered at the German Registry of Clinical Trials (DRKS, URL, registration ID DRKS00030874).
Appendix G Additional Bootstrapping and Wilcoxon Signed-Rank Test Results
Table 6: Bootstrap rank frequency, win probability, and Wilcoxon signed-rank $p$-values for both tasks (F1-score).
Task Center Team Rank Freq. Win Prob. Wilcox $p$
1 2 3 Cam Elb San Cam Elb San
1 4 Cam 0.0000 0.2491 0.7509 NaN 0.2481 0.0000 NaN 0.0000 0.0000
1 4 Elb 0.0077 0.7442 0.2481 0.7509 NaN 0.0077 0.0000 NaN 0.0000
1 4 San 0.9923 0.0077 0.0000 1.0000 0.9923 NaN 0.0000 0.0000 NaN
2 Avg Cam 0.7192 0.1994 0.0814 NaN 0.8501 0.7876 NaN 0.0000 0.0000
2 Avg Elb 0.1023 0.3503 0.5474 0.1499 NaN 0.4050 0.0000 NaN $5.04\times 10^{-106}$
2 Avg San 0.1786 0.4502 0.3712 0.2123 0.5950 NaN 0.0000 $5.04\times 10^{-106}$ NaN
Table 7: Bootstrap rank frequency, win probability, and Wilcoxon signed-rank $p$-values for both tasks (EC).
Task Center Team Rank Freq. Win Prob. Wilcox $p$
1 2 3 Cam Elb San Cam Elb San
1 4 Cam 0.0000 0.0000 1.0000 NaN 0.0000 0.0000 NaN 0.0000 0.0000
1 4 Elb 0.0027 0.9973 0.0000 1.0000 NaN 0.0021 0.0000 NaN 0.0000
1 4 San 0.9979 0.0021 0.0000 1.0000 0.9973 NaN 0.0000 0.0000 NaN
2 Avg Cam 0.2781 0.2955 0.4264 NaN 0.4661 0.3791 NaN $3.13\times 10^{-13}$ $1.15\times 10^{-177}$
2 Avg Elb 0.2855 0.3646 0.3499 0.5307 NaN 0.3991 $3.13\times 10^{-13}$ NaN $1.17\times 10^{-128}$
2 Avg San 0.4412 0.3394 0.2194 0.6176 0.5983 NaN $15\times 10^{-177} 1.1$$17\times 10^{-128}$ NaN
Appendix H Credit Authorship Contribution Statement
Max Kirchner: Conceptualization, Challenge Organization, Methodology, Software, Validation, Formal Analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review and Editing, Visualization, Project Administration Hanna Hoffmann: Conceptualization, Software, Writing - Review and Editing Alexander C. Jenke: Challenge Organization, Software, Writing - Review and Editing Oliver L. Saldanha: Conceptualization, Challenge Organization, Software, Investigation, Data Curation, Writing - Review and Editing Kevin Pfeiffer: Software, Data Curation, Writing - Review and Editing Kanjo Weam: Data Curation, Writing - Review and Editing Julia Alekseenko: Methodology, Software, Investigation, Visualization, Writing - Review and Editing Claas de Boer: Methodology, Software, Investigation, Writing - Review and Editing Santhi Raj Kolamuri: Methodology, Software, Investigation, Writing - Review and Editing Lorenzo Mazza: Methodology, Software, Investigation, Writing - Review and Editing Nicolas Padoy: Challenge Team Supervision, Writing - Review and Editing Sophia Bano: Challenge Organization, Supervision, Writing - Review and Editing Annika Reinke: Conceptualization, Challenge Organization, Supervision, Writing - Review and Editing Lena Maier-Hein: Challenge Organization, Supervision, Writing - Review and Editing Danail Stoyanov: Challenge Organization, Supervision, Writing - Review and Editing Jakob N. Kather: Conceptualization, Challenge Organization, Writing - Review and Editing, Supervision, Project administration, Funding Acquisition Fiona R. Kolbinger: Conceptualization, Challenge Organization, Investigation, Data Curation, Writing - Review and Editing, Supervision, Project Administration, Funding Acquisition Sebastian Bodenstedt: Conceptualization, Challenge Organization, Software, Investigation, Data Curation, Writing - Review and Editing, Supervision, Project Administration, Funding Acquisition, Challenge Team Supervision Stefanie Speidel: Conceptualization, Challenge Organization, Writing - Review and Editing, Supervision, Project Administration, Funding Acquisition, Challenge Team Supervision

### v2

@@
 Abstract
-Purpose: The FedSurg challenge was designed to benchmark the state of the art in federated learning for surgical video classification.
-Its goal was to assess how well current methods generalize to unseen clinical centers and adapt through local fine-tuning while enabling collaborative model development without sharing patient data.
-Methods: Participants developed strategies to classify inflammation stages in appendicitis using a preliminary version of the multi-center Appendix300 video dataset.
-The challenge evaluated two tasks: generalization to an unseen center and center-specific adaptation after fine-tuning.
-Submitted approaches included foundation models with linear probing, metric learning with triplet loss, and various FL aggregation schemes (FedAvg, FedMedian, FedSAM).
-Performance was assessed using F1-score and Expected Cost, with ranking robustness evaluated via bootstrapping and statistical testing.
-Results: In the generalization task, performance across centers was limited.
-In the adaptation task, all teams improved after fine-tuning, though ranking stability was low.
-The ViViT-based submission achieved the strongest overall performance.
-The challenge highlighted limitations in generalization, sensitivity to class imbalance, and difficulties in hyperparameter tuning in decentralized training, while spatiotemporal modeling and context-aware preprocessing emerged as promising strategies.
-Conclusion: The FedSurg Challenge establishes the first benchmark for evaluating FL strategies in surgical video classification.
-Findings highlight the trade-off between local personalization and global robustness, and underscore the importance of architecture choice, preprocessing, and loss design.
-This benchmarking offers a reference point for future development of imbalance-aware, adaptive, and robust FL methods in clinical surgical AI.
+Developing generalizable surgical AI requires multi-institutional data, yet patient privacy constraints preclude direct data sharing, making Federated Learning (FL) a natural candidate solution.
+The application of FL to complex, spatiotemporal surgical video data remains largely unbenchmarked.
+We present the FedSurg Challenge, the first international benchmarking initiative dedicated to FL in surgical vision, evaluated as a proof-of-concept on a multi-center laparoscopic appendectomy dataset (preliminary subset of Appendix300).
+Three submissions were evaluated on generalization to an unseen center and center-specific adaptation.
+Centralized and Swarm Learning baselines isolate the contributions of task difficulty and decentralization to observed performance.
+Even with all data pooled centrally, the task achieved only 26.31% F1-score on the unseen center, while decentralized training introduced an additional, separable performance penalty.
+Temporal modeling emerges as the dominant architectural factor: video-level spatiotemporal models consistently outperformed frame-level approaches regardless of aggregation strategy.
+Naive local fine-tuning leads to classifier collapse on imbalanced local data; structured personalized FL with parameter-efficient fine-tuning represents a more principled path toward center-specific adaptation.
+By characterizing current FL limitations through rigorous statistical analysis, this work establishes a methodological reference point for robust, privacy-preserving AI systems in surgical video analysis.
 keywords:
@@
 Findings on generalization vs. personalization in federated learning
-Demonstrated strengths and limitations of federated surgical AI
+Temporal modeling and personalized FL as key methodological directions
 1 Introduction
@@
 As the first FL challenge of its kind in SDS, FedSurg pioneers the use of a preliminary version of the Appendix300 dataset as a foundational platform for algorithm development and validation.
-A separate benchmarking paper has been published for the complete dataset [22].
+A separate study using the complete dataset is published [42].
 In this paper, we report the challenge design, results, and findings of FedSurg according to the transparent reporting of biomedical image analysis challenges (BIAS) guidelines [34].
@@
 This challenge’s main contribution is its patient-level prediction task for appendicitis staging, a transferable innovation for AI in surgery.
-By establishing a framework for standardized, objective assessment of inflammation severity, this approach serves as a powerful proof-of-concept for addressing more critical clinical needs.
-These include advancing intraoperative assistance for surgical quality control, and providing a foundation for developing AI tools for accurate diagnosis, real-time surgical support, and effective prognosis for a variety of surgical conditions.
+This challenge serves as a proof-of-concept for applying FL to patient-level surgical video analysis, with methodological insights intended to transfer to future tasks with broader clinical relevance, such as intraoperative decision support or quality control across diverse surgical procedures.
 Participants were tasked with developing FL algorithms to classify appendicitis stages in laparoscopic videos, structured around two core tasks:
@@
 2.3 Challenge Dataset
-The challenge dataset is a preliminary subset of the Appendix300 collection, comprising frames from 223 full-length recordings of laparoscopic appendectomies.
+The challenge dataset is a preliminary subset of the Appendix300 collection and comprises frames from 223 full-length recordings of laparoscopic appendectomies.
 Beyond image frames and laparoscopic grading, the finalized Appendix300 dataset is enriched with detailed histopathological findings and patient anamnesis data [24].
@@
 An overview of the video distribution across centers is shown in Table 1.
-Table 1: Data Distribution Across Participating Centers.
+Table 1: Data Distribution Across Participating Centers. Every video is standardized to a 100-second clip, centered on the timestamp of full appendiceal visibility prior to surgical preparation.
 Center Training Videos Testing Videos Total
@@
 Label distributions for (a) the training dataset and (b) the test dataset across the four centers.
+In the training set visualization, the darker segments represent the publicly available subset for participant development, while the lighter segments show the complete dataset used for final federated training.
 The plots highlight notable inter-center variability and class imbalance.
-In the training set visualization, the darker segments represent the publicly available subset for participant development, while the lighter segments show the complete dataset used for final federated training.
-The exact data distribution was unknown for the participants.
 To discourage centralized model development and preserve the integrity of the federated setting, the training data was partitioned into public and private subsets.
@@
 Final submissions were executed centrally by the organizers, who conducted full federated training and evaluation across all centers under secure and standardized conditions to ensure reproducibility and fairness.
-Since the challenge dataset constitutes a preliminary subset of the Appendix300 dataset, we provide a CSV file specifying which samples were used in the challenge here [^1] Available upon acceptance or on request..
+Since the challenge dataset constitutes a preliminary subset of the Appendix300 dataset, we provide a CSV file specifying which samples were used in the challenge here [^1] Available at https://nextcloud.tso.ukdd.de/s/XppWkRnWpLGwj5F Password for Review: FedSurg24_data!.
 In addition, two videos are released that were excluded from the final dataset because they were either shorter than 100 seconds or did not contain a clearly visible appendix in the center.
@@
 The details of our evaluation procedure are described below.
-For Task 1 (generalization), we computed the F1-score and the EC metric for all test cases.
+For Task 1 (generalization), we computed the F1-score and the EC for all test cases.
 Separate rankings were assigned based on each metric, and a team’s overall rank for Task 1 was derived by averaging these two ranks.
@@
 Kolamuri, who submitted independently as Team Santhi; Lorenzo Mazana and Claas de Boer from the Translational Surgical Oncology group at the National Center for Tumor Diseases, submitting as Team Elbflorenz; and Julia Alekseenko and Nicolas Padoy from the CAMMA research group at IHU Strasbourg, submitting as Team Camma.
+Furthermore, we provide centralized baseline and a federated Swarm Learning (SL) baseline.
 3.1 Participating Teams and Methods
@@
 Camma observed that prototype-based inference worked best for Centers 2 and 3, while per-sample comparison was more effective for Center 1, allowing the approach to flexibly adapt to center-specific data distributions.
-This metric learning paradigm is particularly well-suited for FL, as it focuses on learning a robust embedding space rather than simply aggregating classifier weights, which can be highly sensitive to the non-IID label distributions across centers.
+This metric learning paradigm is particularly well-suited for FL, as it focuses on learning a robust embedding space rather than simply aggregating classifier weights, which can be highly sensitive to the non-independent-and-identically-distributed (non-IID) label distributions across centers.
 To address domain heterogeneity, the model incorporates Switchable Normalization [30], which combines batch, instance, and layer normalization through softmax gating.
@@
 Overall, Team Camma’s method demonstrates how metric learning, adaptive normalization, and robust aggregation can be combined for scalable, personalized FL in surgical video classification.
+3.2 Baselines
+In addition to the three participating teams, we provide two baselines for comparison: a centralized baseline and a federated baseline. Both baselines are based on the SurgTempoNet method [42].
+SurgTempoNet employs a spatial-temporal architecture designed to predict patient-level clinical outcomes directly from laparoscopic videos.
+The model utilizes a ConvNeXt-Base [29] backbone, pre-trained on ImageNet-22k [41], to extract high-level feature representations from individual frames.
+In its temporal configuration, these features are fed into a single-layer Long Short-Term Memory (LSTM) [18] network with a hidden state size of $160$.
+This allows the model to aggregate sequential information across the entire surgical procedure, mapping a sequence of frames to a single global label through a final linear classification layer.
+To manage the high dimensionality of surgical footage, the pipeline uses a weakly supervised approach with sparse frame sampling.
+Videos are processed at a rate of one frame per second, and frames are resized to a resolution of $480\times 270$ pixels.
+This sampling strategy reduces the computational burden of processing every raw video frame.
+Both centralized and federated training is conducted using the AdamW optimizer with a learning rate of $1.0\times 10^{-3}$.
+To address the inherent class imbalance in surgical datasets, the model utilizes Weighted Cross-Entropy Loss, where weights are inversely proportional to the frequency of each disease stage in the training set.
+Furthermore, a batch accumulation strategy is implemented, where gradients are computed over mini-batches of $64$ clips, but weights are updated only every three steps, effectively stabilizing the training process for video sequences.
+The defining characteristic of the federated baseline is its integration into an SL framework [47, 17].
+Unlike centralized training or standard FL, local models are trained on site-specific data, and only the model weights are synchronized via a peer-to-peer network managed by a SwarmCallback function.
+This synchronization occurs at a fixed frequency (e.g., every $100$ batches) and requires a minimum number of active peers to proceed.
+Furthermore, to probe whether parameter-efficient fine-tuning (PEFT) constitutes a viable adaptation strategy in this data regime, two additional preliminary models are included.
+Both apply Low Rank Adaptation (LoRA) [19] (rank 8, alpha 16) to the attention projections of a frozen backbone, combined with a shared three-layer MLP classification head and FedAvg aggregation over 25 rounds; all remaining hyperparameters (ASAM optimizer, cosine annealing, four local epochs per round) are held constant.
+[…]
 3.3 Scores and Rankings
 The performance of the submitted federated models was evaluated in both tasks of the FedSurg Challenge: generalization to an unseen center (Task 1) and center-specific adaptation (Task 2).
-Performance was measured using Expected Cost (EC) and F1-score.
+Performance was measured using EC and F1-score.
 Rank stability was assessed via 10,000-iteration bootstrapping, followed by the Wilcoxon signed-rank test for statistical significance.
-A comprehensive summary of results—including key metrics, confusion matrices, and rank stability—is provided in Tables 3 and 4 and Figures 4, 5, 6, and 7, with further details in Appendix G.
+Additionally, the baseline methods are reported separately in Section 3.4 as contextual references and are not part of the official challenge ranking.
+Overall, a comprehensive summary of results, including key metrics, confusion matrices, and rank stability, is provided in Tables 3 and 4 and Figures 4, 5, 6, and 7, with further details in G.
 3.3.1 Task 1: Generalization to an Unseen Center
@@
 The poor performance of this trivial classifier is supported by bootstrapping analysis (Figure 5) and statistical significance testing (Table 3).
-Across 10,000 repetitions, Team Camma consistently ranked as the lowest-performing team in 100% of cases for the EC-Score and 75.09% for the F1-score (Figure 6).
+Across 10,000 repetitions, Team Camma consistently ranked as the lowest-performing team in 100% of cases for the EC and 75.09% for the F1-score (Figure 6).
 Team Elbflorenz also demonstrated limited performance, predominantly predicting class 2 across inputs (Figure 4).
@@
 The confusion matrices in Figure 7 reveal predictions largely concentrated along the diagonal, suggesting good class differentiation.
-In Center 3, however, the local model—similar to Team Elbflorenz—showed limited discriminatory power, with a tendency to predict classes 2, 3, or 4 rather than providing balanced classification.
+In Center 3, however, the local model, similar to Team Elbflorenz, showed limited discriminatory power, with a tendency to predict classes 2, 3, or 4 rather than providing balanced classification.
 Bootstrapping analysis (Figure 5) confirms consistent performance across centers, but the final ranking remained unstable, as indicated by the wide 95% confidence intervals in Figure 6.
@@
 Santhi 1 1 1 1 2 1 1
+3.4 Further Analysis
+Challenge participant results are compared against organizer-provided baselines comprising centralized training, SL, and parameter-efficient fine-tuning (PEFT) variants to disentangle intrinsic task difficulty from federated overhead.
+Communication efficiency across all evaluated approaches is analyzed subsequently.
+Ranking and evaluation methodology are described in Section 2.4.1.
+3.4.1 Baseline comparison
+Table 6: Absolute raw test performance comparison including baselines.
+EC: lower is better; F1: higher is better.
+Baseline rows in the lower half of the table are contextual references and are not part of the official challenge ranking.
+Method Task 1 EC Task 1 F1 Task 2 Avg EC Task 2 Avg F1
+Camma 57.24% 4.76% 21.79% 18.91%
+Elbflorenz 24.14% 7.83% 21.35% 13.14%
+Santhi 12.41% 23.03% 20.44% 15.40%
+Cen 10.34% 26.31% 19.82% 23.51%
+SWARM 13.10% 19.00% 20.41% 17.77%
+EndoViT+LoRA 17.24% 9.24% 20.13% 8.70%
+ViViT+LoRA 22.76% 15.93% 17.60% 21.47%
+Table 7: Center-wise Task 2 baseline results. EC: lower is better; F1: higher is better.
+Baseline Center 1 Center 2 Center 3
+EC F1 EC F1 EC F1
+Cen 24.00% 11.43% 20.00% 26.43% 15.45% 32.67%
+SWARM 28.00% 10.32% 17.78% 20.95% 15.45% 22.03%
+EndoViT+LoRA 17.78% 10.32% 24.44% 6.06% 18.18% 9.68%
+ViViT+LoRA 22.22% 12.22% 13.33% 22.62% 17.27% 29.57%
+To provide a comprehensive benchmark, we introduced four baselines.
+A centralized (Cen) baseline, representing the upper bound of performance by training on pooled data from all centers for Task 1 and on each center’s data separately for Task 2, and an SL baseline, applying the same architecture in a decentralized peer-to-peer (P2P) setup, together quantify the cost of decentralization.
+Two additional PEFT baselines, EndoViT+LoRA and ViViT+LoRA, probe whether parameter-efficient adaptation can serve as a practical alternative to full local fine-tuning.
+Table 6 summarizes absolute raw test performance across both tasks, while Table 7 reports center-wise Task 2 results for all baseline models.
+Comparative Performance
+The centralized baseline outperformed all decentralized models in both tasks, achieving the top rank in both F1-score and EC metrics.
+In Task 1 (Generalization), the centralized baseline reached an F1-score of $26.31\%$ and an EC of $10.34\%$, followed by Santhi and the SL baseline (compare Figure 9).
+Camma and Elbflorenz recorded the lowest Task 1 performance among all methods, with F1-scores of 4.76% and 7.83% respectively.
+In Task 2 (Adaptation), the performance gap narrowed, reflecting the challenge of center-specific fine-tuning.
+While the centralized model maintained the lead (Avg.
+F1: $23.51\%$), the SL baseline demonstrated strong adaptation across the three centers, securing the second overall rank for this task besides Santhi’s method (compare Table 11).
+Camma and Elbflorenz both improved substantially in Task 2 relative to their Task 1 performance, yet neither reached the performance level of the centralized or SL baselines.
+The two PEFT baselines reveal a clear separation by architecture.
+EndoViT+LoRA achieved 9.24% F1 on Task 1 and only 8.70% average F1 on Task 2, performing comparably to Elbflorenz, the frame-level method it is derived from, but failing to surpass it.
+Inspection of per-class predictions confirms near-complete collapse to the majority class, indicating that LoRA adaptation of a frame-level backbone is insufficient to overcome the high intra-video visual variance at an unseen center.
+ViViT+LoRA, by contrast, achieved 15.93% F1 on Task 1 and 21.47% average F1 on Task 2, ranking second among all methods in Task 2.
+Notably, at Center 3, ViViT+LoRA reached 29.57% F1, the highest Task 2 F1 among all non-centralized methods and approaching the centralized baseline (32.67%).
+Performance at the smaller centers remained limited (Center 1: 12.22%, Center 2: 22.62%), consistent with the quantity-skew effect described in Section 4.3: local adaptation via PEFT benefits from sufficient local samples and is most effective at the largest node.
+Rank Stability and Statistical Significance
+To address the stability of these findings beyond point estimates, we conducted a stability analysis using 10,000-iteration bootstrapping.
+This process quantifies the probability of each team achieving a specific rank under different data distributions, as visualized in Figure 9.
+As shown in Figure 9(a) and (b), the rankings for Task 1 were highly stable with a tendency to solve the ties in the ranking.
+For the F1-score, the centralized baseline retained the 1st rank in $67.0\%$ of iterations, while SWARM maintained the 3rd rank in $72.7\%$ of cases.
+In contrast, Task 2 exhibited higher volatility (Figure 9(c) and (d)), particularly in the EC metric.
+For instance, Camma’s rank for Task 2 EC was distributed across all five positions (retaining the 5th rank in only $31.9\%$ of iterations), reflecting a higher sensitivity to the specific data composition of the adaptation centers.
+The statistical significance of these rankings was validated using the Wilcoxon signed-rank test.
+In Task 1, all pairwise comparisons for both metrics yielded $p<0.001$, indicating that the performance tiers between the baselines and participants are statistically distinct.
+In Task 2, while most differences remained highly significant ($p\approx 0$), the comparison between Santhi and SWARM for the EC metric yielded a $p$-value of $0.745$.
+This suggests that their adaptation performance regarding surgical count estimation was statistically indistinguishable, despite their divergence in F1-score ($p<0.001$).
+The variability of these outcomes is further illustrated in Figure 8, where error bars represent the standard deviation across bootstrap replicates.
+The narrower distributions for Task 1 compared to the wider variance in Task 2 underscore the inherent difficulty of the adaptation task, where local center characteristics significantly influence model stability.
+Figure 8: Bootstrapped Performance Results with Baselines.
+Visualization of the performance results with standard deviation as error bars for all teams and tasks after bootstrapping with 10,000 repetitions.
+The plot illustrates the variability and stability of the outcomes across different centers.
+Figure 9: Ranking Stability.
+Bootstrapped ranking distributions for each metric and task, based on 10,000 bootstrap iterations.
+Circle size indicates the percentage of times a team’s model achieved a specific rank across samples.
+Black crosses show median ranks, and black lines denote the 95% bootstrap confidence intervals.
+Subfigures (a) and (b) correspond to Task 1 (generalization ability) with metrics EC and F1-score, respectively, while (c) and (d) represent Task 2 (adaptation ability) using the same metrics.
+3.4.2 Communication efficiency analysis
+Communication cost is a critical metric in decentralized machine learning because the frequent transmission of high-dimensional model parameters, weights, or gradients over limited or unstable network bandwidth can become a significant bottleneck [36].
+In many medical AI applications, the time required for network synchronization often dwarfs the time spent on local GPU computation.
+Consequently, optimizing these costs is essential for making collaborative learning scalable and practical for real-world clinical scenarios.
+In a traditional FL architecture, the total communication costs can be calculated using the following formula:
+$C_{\text{total}}=2\times K\times T\times|W|$ (5)
+Where:
+  • 1.
+    $C_{\text{total}}$ is the total communication cost (data transferred).
+  • 2.
+    The factor of $2$ accounts for the bi-directional transfer (upload and download) of model parameters between the client and the central server in each round.
+  • 3.
+    $K$ is the number of participating clients (or nodes) per round.
+  • 4.
+    $T$ is the total number of communication rounds.
+  • 5.
+    $|W|$ represents the size of the model parameters or weights.
+However, for SL, the communication cost is calculated differently due to its peer-to-peer (P2P) nature.
+Since a central server does not exist, nodes synchronize weights directly through a decentralized mesh.
+In a leader-based synchronization (as utilized in the SurgTempoNet framework), one node is elected to aggregate weights from its peers.
+Consequently, the communication cost formula is adjusted to reflect the K1K-1 active connections:
+$C_{\text{total}}=2\times(K-1)\times S\times|W|$ (6)
+Where $S$ represents the total number of synchronization events across the entire training duration.
+In the SurgTempoNet implementation, synchronization is governed by a frequency parameter rather than fixed rounds.
+The total number of synchronization events $S$ for a node is defined as:
+$S=\sum_{n=i}^{K}\left\lfloor\frac{E\times B_{i}}{f_{\text{sync}}}\right\rfloor+1$ (7)
+Where:
+  • 1.
+    $E$ is the number of local epochs performed.
+  • 2.
+    $B_{i}$ is the number of batches contained in each epoch for center $i$.
+  • 3.
+    $f_{\text{sync}}$ is the synchronization frequency (e.g., every 100 batches).
+  • 4.
+    The +1+1 accounts for the initial synchronization at the start of training.
+By utilizing 2(K1)2(K-1), the model accounts for the fact that the aggregating node does not incur external network traffic for its own local weights.
+This provides a more rigorous representation of the physical network traffic generated during the decentralized training of surgical video classifiers across heterogeneous hospital sites.
+Table 8 reveals again significant variance in both architectural complexity and the resulting communication overhead during the FL process.
+Model Architecture and Paramerization
+The models utilized by Santhi, Elbflorenz, and our SL Baseline share a similar scale, ranging around 88M parameters.
+In contrast, team Camma implemented a significantly more compact architecture.
+With only 24M parameters, Camma’s model is approximately 73 % smaller than the other methods.
+All models utilize 32-bit floating-point precision, resulting in a model size of approximately 340 MB for the larger architectures and around 92 MB for Camma’s compact model.
+Synchronization Events and Total Communication Costs
+As discussed, the total communication cost is a function of both the model size and the frequency of synchronisation events.
+The SL Baseline, with its high synchronization frequency (100 events), incurs the highest communication cost, totaling approximately 133.14 GB across all clients and rounds, despite saving communication compared to the other methods by using a P2P approach.
+Furthermore, the impact of synchronization frequency is most evident when comparing Santhi and Elbflorenz.
+Team Elbflorenz, with 50 synchronization events, results in a total cost of around 97.47 GB.
+Team Santhi, with only 20 synchronization events, significantly reduces the communication overhead to approximately 39.63 GB.
+Finally, team Camma, with the smallest model and lowest synchronization frequency (10 events), achieves the lowest communication cost of about 5.38 GB.
+In summary, team Camma emerged as the most communication-efficient team by a wide margin.
+By combining a reduced parameter count with a highly conservative synchronization strategy (10 sync events), they achieved a total communication cost of only 5.38 GB.
+Compared to the SL Baseline, this represents a $95.9\%$ reduction in network traffic, highlighting the efficacy of aggressive model compression and sparse synchronization in federated environments.
+In the current implementation, both PEFT baselines transmitted full model snapshots during aggregation, incurring communication costs comparable to full-model FL methods.
+Since only the LoRA adapter weights differ between federation rounds ($\sim$940–969K parameters, $\sim$3.8 MB per model), a dedicated LoRA-only synchronization strategy would reduce per-round transmission to less than 2% of the costs reported in Table 8.
+This efficiency advantage was not exploited here, but represents a compelling practical benefit of PEFT-based FL for future implementations.
+Table 8: Summary of architectural parameters and cumulative synchronization data costs per team.
+Team SL Baseline Santhi Elbflorenz Camma
+Parameters 89,351,470 88,651,014 87,213,038 24,086,014
+Data Type float32 (4 bytes/param) float32 (4 bytes/param) float32 (4 bytes/param) float32 (4 bytes/param)
+Model size 357,405,880 B (340.85 MB) 354,604,056 B (338.18 MB) 348,852,152 B (332.69 MB) 96,344,056 B (91.88 MB)
+Clients 3 3 3 3
+Sync Events 100 20 50 10
+Total Costs 142,962,352,000 B (133.14 GB) 42,552,486,720 B (39.63 GB) 104,655,645,600 B (97.47 GB) 5,780,643,360 B (5.38 GB)
 4 Discussion
-4.1 Task Wise performance
-The FedSurg Challenge establishes the first benchmark for evaluating Federated Learning (FL) strategies in the surgical video classification of appendicitis grades, utilizing a heterogeneous, multi-institutional dataset.
-Three complete submissions were received and evaluated on a preliminary subset of the Appendix300 dataset.
-Task 1, which focused on generalization to an unseen center, immediately revealed the limited generalization abilities of current FL models.
-This difficulty stems directly from significant data heterogeneity across the participating institutions.
-For instance, images originated from a wide array of surgical video systems, including Arthrex Synergy UHD4, Storz image1 s, B.
-Braun Aesculap EinsteinVision, and Richard Wolf Endocam 4k, leading to variations in lighting and resolution [21].
-For example, Center 1 provided 4k images, while Center 4 provided cropped images in Full HD resolution.
-In addition to this feature skew, a label distribution skew existed between the different centers (Figure 2).
-In this challenging environment, models capable of temporal modeling, like the ViViT-based submission, achieved the strongest but still weak results.
-A slight tendency also emerged showing that approaches focused on the loss landscape, like FedSAM, performed slightly better than a FedMedian aggregation of a triplet margin loss and cosine similarity model.
-This aligns with the conclusions of Foret et al.
-[39] and Caldarola et al.
-[38] that the quest for flat minima, which is incorporated into FedSAM, can lead to better generalization ability.
-This difficulty was not unique to the challenge submissions; a benchmark study on the Appendix300 dataset [22] using a leave-one-out cross-validation design similarly confirmed the profound challenge of inter-center generalization.
-Task 2, which involved center-specific adaptation after fine-tuning, led to only marginally better results across all teams.
-This marginal improvement, however, came at the cost of rank instability, highlighting that even minor gains in local adaptation can compromise the robustness of a model’s comparative performance.
-The limited improvement can be attributed to several factors inherent to the federated setup.
-Due to the difficult knowledge of the complete data distribution and amounts, hyperparameter tuning was challenging, which could lead to local models overfitting to their local training data and, therefore, still failing to achieve high accuracy.
-Furthermore, the knowledge from other centers within the global model can negatively influence client performance.
-At each training round, the client initializes its weights from the global model, which represents an aggregated compromise of the knowledge learned across all participating centers.
-The poor local performance could stem from conflicting updates that pull the model away from what is optimal for its own center’s data.
-Ultimately, these findings illustrate the central challenge of balancing generalization and adaptation, alongside the overarching difficulty of data heterogeneity in federated setups.
-4.2 Stability assessment
-The bootstrapped rank frequency analysis yielded distinct outcomes for Task 1 and Task 2.
-For Task 1, the analysis indicated stable leader board rankings, with the predictions from submitted models being statistically distinct.
-A central finding was the critical role of temporal information; the highest-ranking submission utilized a video-based model, whereas models predicated on static frames exhibited the poorest performance.
-This aligns with the findings from the Appendix300 benchmarking paper [22], where the model trained only on middle frames performed poorly in comparison to the models trained on video models with a temporal horizon.
-Furthermore, using a higher frame rate up to 1 frame per second (fps) resulted into better results.
-In stark contrast, the rankings for Task 2 demonstrated pronounced instability.
-No submission maintained its rank on the EC metric in more than 50% of the bootstrap replicates, indicating that the leader board outcomes were not robust.
-This instability suggests that leader boards for federated adaptation tasks should be interpreted with extreme caution, as top-ranking methods may not be reliably superior but merely fortunate in a specific data configuration.
-4.3 Systemic Limitations and Data-Related Challenges in Federated Surgical AI
-Several systemic issues emerged.
-FL does not resolve underlying dataset problems such as class imbalance and can amplify heterogeneity effects.
-Rare inflammation stages were underrepresented, leading to poor classification across all teams.
-Personalized losses and fine-tuning improved performance within centers (e.g., Team Camma) but compromised generalization to unseen centers, underscoring a fundamental trade-off in FL.
-This mirrors our findings from the primary tasks, where the unstable gains in center-specific adaptation (Task 2) failed to translate into the robust, generalizable models needed for unseen centers (Task 1).
-Similarly, design choices impacted outcomes: lightweight frozen foundation models provided stability across centers, whereas more adaptive approaches (e.g., Siamese networks with triplet loss) improved local performance but struggled with generalization under distribution shifts.
-Furthermore, hyperparameter tuning proved difficult without access to the true data distribution, reflecting real-world FL constraints.
-The unknown class distribution contributed to lower-than-expected performance.
-Moreover, human experts showed only moderate agreement on inflammation grading ($\kappa_{weighted}=0.62$ [21]), underscoring the inherent ambiguity of the classification task.
-4.4 Future directions
-Future work should address identified limitations.
-Imbalance-aware strategies, such as reweighting, resampling, or federated focal loss, are needed to improve performance on clinically critical minority classes.
-Context-aware preprocessing and targeted frame sampling consistently outperformed uniform strategies and should be further explored.
-Methodological advances such as federated domain adaptation, personalized aggregation, and uncertainty-aware inference may improve robustness and safety.
-Finally, self-supervised pretraining on large-scale surgical video datasets could yield more transferable representations without compromising data privacy.
-4.5 Challenge insights
-Despite 24 registered teams, only three complete submissions were ultimately received.
-This discrepancy reflects both the niche character of FL within surgical AI and the complexity of the challenge setup.
-Unlike conventional challenges, FedSurg required fully encapsulated end-to-end submissions in Docker, with limited access to data and restricted debugging capabilities.
-These hurdles highlight practical barriers to broader participation and point to the technical demands of real-world FL applications.
+The FedSurg challenge should be understood as a proof-of-concept benchmark: the appendectomy task was selected for its methodological tractability rather than for its immediate clinical impact.
+It offers a standardized surgical workflow, multi-center data availability, and an established grading system.
+The following discussion focuses on methodological insights transferable to future FL applications on surgical video, structured around two central questions: first, what architectural and algorithmic choices best support global generalization to unseen centers; and second, whether and how local adaptation can be reliably achieved without sacrificing the globally learned representation.
+4.1 The Role of Temporal Modeling
+A key finding of this challenge is that temporal modeling strategy is the dominant performance factor, outweighing the choice of aggregation scheme or communication frequency.
+Frame-level approaches, whether operating through majority voting (Team Elbflorenz) or embedding-based prototype comparison (Team Camma), are inherently limited by the high intra-video visual variance present in laparoscopic recordings.
+Differences in laparoscopic hardware, lighting conditions, and patient anatomy render individual frames unreliable predictors of patient-level inflammation grade.
+The metric learning approach of Team Camma illustrates this most clearly: while it achieved competitive performance at centers whose visual characteristics were represented in training, it failed to map embeddings of the unseen center to established training prototypes, revealing the brittleness of a feature space built on individual frames.
+Video-level temporal models, by contrast, capture the progression of visual features across the surgical sequence, providing contextual information no single frame can supply.
+The ViViT-based submission of Team Santhi, processing 32 frames jointly, consistently achieved the strongest generalization to the unseen center in Task 1, outperforming both frame-level approaches despite using fewer communication resources.
+This finding is consistent with prior work on decentralized surgical video analysis using the same dataset [42], in which temporal architectures similarly outperformed frame-level methods across patient-level prediction tasks.
+With respect to the generalization vs.
+adaptation trade-off, temporal modeling also offered a partial advantage in Task 2: Team Santhi’s frozen ViViT backbone preserved global spatiotemporal representations during local fine-tuning, limiting the degree to which center-specific updates could corrupt the globally learned feature space.
+This suggests that architecture choice is not only a generalization strategy but also a safeguard against the instability of local adaptation.
+This point is revisited in Section 4.4.
+4.2 Dataset Scale and Decentralization as Compounding Factors
+The generally low absolute scores across all submissions raise an important question: do these results reflect fundamental limitations of FL, or primarily the difficulty of the task itself?
+The centralized baseline provides a direct answer.
+Even when pooling all multi-institutional data onto a single server, thereby removing all federated constraints, the centralized model achieved only 26.31% F1-score on the unseen center.
+This establishes that the underlying task of patient-level appendicitis grading from surgical video is intrinsically difficult, regardless of the learning paradigm employed, due to the high visual variance across centers and the inherent ambiguity of fine-grained clinical grading.
+Decentralization acts as an additional, separable bottleneck.
+The Swarm Learning baseline, which retains the same architecture as the centralized model but distributes training across centers, achieved 19.00% F1 in Task 1.
+This measurable degradation is attributable specifically to decentralized optimization across non-IID data nodes, rather than to any architectural deficiency.
+This gap quantifies what may be termed a decentralization penalty: the cost in predictive performance of preserving data privacy through distributed training.
+Together, the two baselines confirm that observed performance reflects a compounding of two distinct factors, and that algorithmic advances in FL aggregation must overcome both.
+An additional consideration is the scale of the dataset.
+The challenge utilized 153 training videos, which constitute a preliminary subset of the 330-video Appendix300 dataset [24], distributed across four centers.
+At this data regime, FL fragments an already limited training pool, leaving individual nodes with insufficient samples to learn robust local representations.
+Prior work using the full dataset [42] demonstrated improved performance under decentralized training, suggesting that the current performance ceiling is partly a function of data scarcity rather than an inherent ceiling of the FL paradigm.
+Scaling the dataset, in terms of both total videos and number of contributing centers, is therefore a necessary condition for a more definitive evaluation of FL on surgical video.
+4.3 Data Heterogeneity and the Limits of Local Adaptation
+The challenge results highlight that non-IID data remains the central unsolved problem in federated surgical video analysis.
+Distribution shift across centers originates from two compounding sources: hardware-induced feature skew, including differences in laparoscopic systems, resolution, lighting, and cropping, and patient population heterogeneity, including differences in age distribution, disease severity prevalence, and surgical practice between academic and community centers.
+Current aggregation strategies, including FedAvg and FedSAM, partially mitigate but do not resolve these shifts.
+These effects are further amplified by inter-center data quantity skew.
+Center 3 contributed 80 of the 153 training videos, causing global models to disproportionately reflect Center 3’s data distribution.
+The bootstrapped rank stability analysis illustrates the consequence: Task 1 rankings were stable, confirming that temporally-informed models were statistically distinguishable; Task 2 rankings showed pronounced instability, with no submission retaining its rank on the EC metric in more than 50% of bootstrap replicates.
+Local fine-tuning at smaller centers triggered reversion to trivial classifiers, predominantly predicting a single class as a consequence of catastrophic forgetting on heavily skewed local data, inflating nominal F1-scores without reflecting genuine adaptation.
+As discussed in Section 2.4, the EC metric was more robust to this artifact and should be considered the primary indicator of adaptation quality in imbalanced ordinal settings.
+4.4 Toward Personalized Federated Learning
+The results of both tasks point toward the same conclusion: a single globally aggregated model is a compromise that satisfies no center optimally.
+However, naive local fine-tuning, as evaluated in Task 2, does not resolve this, as unconstrained adaptation on skewed local data leads to catastrophic forgetting rather than genuine personalization.
+The path forward lies in structured personalized FL: approaches that separate a shared global representation from center-specific adaptation components, and train both with objectives that explicitly account for local data imbalance.
+For Task 1 generalization, the FedSurg results support the use of a frozen pretrained temporal backbone as the globally federated component.
+By fixing backbone weights across federation rounds, center-specific visual variance cannot propagate into the shared representation, preserving its generalizability to unseen centers.
+Only lightweight task-specific heads need to be aggregated globally, substantially reducing both communication cost and the risk of conflicting updates between heterogeneous nodes.
+For Task 2 adaptation, parameter-efficient fine-tuning (PEFT) methods offer a principled alternative to full local fine-tuning.
+Methods such as LoRA or adapter layers constrain the parameter space available for local updates, limiting the degree to which a small, imbalanced local dataset can destabilize the global representation, while still enabling meaningful center-specific adjustment.
+This approach is particularly relevant given the data regime of this challenge, where local node datasets ranged from 33 to 80 videos.
+Looking further ahead, surgical video foundation models represent a promising direction for initializing the shared backbone [48, 43].
+However, currently available surgical foundation models are trained on limited procedure-specific data, and general-purpose video foundation models produce feature spaces too broad for the fine-grained visual distinctions required in intraoperative grading tasks.
+Recent work combining FL with endoscopic foundation model pretraining [22] suggests a more principled path: federating the pretraining process itself to build privacy-preserving, domain-specific representations from the outset.
+The integration of such approaches with PEFT-based local adaptation remains a hypothesis for future evaluation rather than a validated recommendation.
+The exploratory PEFT baselines included in this work provide initial empirical grounding for these recommendations.
+The EndoViT+LoRA variant, which applies LoRA to a frozen frame-level backbone, collapsed to majority-class prediction (Task 1 F1: 9.24%), confirming that parameter-efficient adaptation cannot compensate for the absence of temporal modeling in this data regime.
+The ViViT+LoRA variant, applying LoRA to a frozen temporal backbone, achieved 15.93% F1 on Task 1 and 29.57% F1 at Center 3 in Task 2, approaching the centralized baseline (32.67%) at the largest local node.
+Performance at smaller centers remained limited (Center 1: 12.22%, Center 2: 22.62%).
+This is consistent with the quantity-skew hypothesis: PEFT-based local adaptation is most effective where sufficient local samples are available to support stable fine-tuning.
+4.5 Implications for Task Design
+Beyond algorithmic choices, the FedSurg results suggest that the design of the clinical prediction task itself should inform the choice of FL strategy.
+This dimension has received limited attention in the literature.
+The consequences of prediction errors differ fundamentally between tasks, and these differences have direct implications for how generalization and adaptation should be balanced.
+Consider two contrasting task types.
+In high-sensitivity settings, for example the detection of malignant lesions in colonoscopy where false negatives carry direct patient harm, global generalization to unseen centers is critical.
+FL strategies that prioritize robustness across non-IID distributions, such as flat-minima optimization approaches like FedSAM, are preferable over those that optimize local performance at the cost of cross-center generalizability.
+The fine-grained 6-class inflammation grading evaluated in this challenge approaches the opposite end of the spectrum: overcalling severity could influence clinical decision-making, placing greater weight on specificity.
+Here, the instability of local adaptation observed in Task 2 is particularly concerning.
+Classifier collapse, the reversion to a trivial classifier predominantly predicting a single class as a consequence of catastrophic forgetting during fine-tuning on imbalanced local data, is not only a performance failure but a risk in high-specificity grading contexts.
+These observations lead to a practical hypothesis for future work: FL strategy selection should be guided not only by dataset characteristics, but by an explicit analysis of the clinical error cost structure of the target task.
+High-sensitivity tasks call for globally robust aggregation; high-specificity tasks require adaptation strategies that prevent classifier collapse, such as the PEFT-based approaches discussed in Section 4.4.
+Systematic evaluation of this hypothesis across diverse surgical AI tasks remains an important open problem.
+4.6 Limitations and Future Directions
+Several limitations of this challenge should be acknowledged.
+Despite 24 registered teams, only three complete submissions were received, reflecting both the niche character of FL within surgical AI and the high technical barriers to participation.
+The dataset, as a preliminary subset of Appendix300 [24], was insufficient in scale for robust federated training, and the simulated FL setup, where all data resided centrally during evaluation, does not capture the full complexity of real-world distributed deployments, including network latency, node availability, and hardware heterogeneity.
+Future work should prioritize scaling the dataset across a broader network of institutions and transitioning from simulated to genuine multi-institutional FL.
+On the algorithmic side, imbalance-aware objectives, such as federated focal loss or client-weighted aggregation, and personalized aggregation methods that account for quantity skew are necessary next steps.
+Self-supervised pretraining on large-scale surgical video collections, ideally in a federated manner [22], offers a promising route toward more transferable representations without compromising data privacy.
 5 Conclusion
-The FedSurg Challenge provides the first benchmark for FL in surgical video classification, highlighting both opportunities and current limitations.
-The findings demonstrate that while spatiotemporal modeling and preprocessing choices can support generalization and adaptation, challenges persist in handling data imbalance, hyperparameter tuning, and stability of results.
-By exposing these trade-offs, FedSurg establishes an important reference point for the surgical AI community and guides future work toward robust, adaptive, and clinically relevant FL methods.
+The FedSurg Challenge establishes the first benchmark for FL in surgical video classification, evaluated here as a proof-of-concept using a multi-center dataset of laparoscopic appendectomies.
+The results reveal that temporal modeling is the dominant architectural factor for cross-center generalization, while decentralized training introduces a measurable performance penalty on top of the intrinsic difficulty of the task.
+Local fine-tuning without imbalance-aware objectives leads to classifier collapse rather than genuine personalization, underscoring the need for structured personalized FL with parameter-efficient adaptation.
+By characterizing these failure modes through rigorous statistical analysis, this work provides a methodological reference point for the design of robust, privacy-preserving AI systems for surgical video analysis and motivates future benchmarking on tasks with broader clinical scope.
 Acknowledgments
@@
 Declaration of generative AI and AI-assisted technologies in the writing process.
-During the preparation of this work, the author(s) used ChatGPT, Gemini 2.5 Pro and DeepL Write in order to improve the readability and language of the manuscript.
+During the preparation of this work, the author(s) used ChatGPT, Gemini 2.5 Pro, Claude Code, and DeepL Write in order to improve the readability and language of the manuscript.
 After using these tools/services, the author(s) reviewed and edited the content as needed and take full responsibility for the content of the published article.
@@
 The challenge dataset is a preliminary subset of the publicly available Appendix300 dataset [24].
-A CSV file detailing the samples used in the challenge is available here  [^2] Available upon acceptance or on request., along with two videos that were excluded from the final dataset but used in the challenge for reproducibility.
+A CSV file detailing the samples used in the challenge is available here  [^2] Available at https://nextcloud.tso.ukdd.de/s/XppWkRnWpLGwj5F Password for Review: FedSurg24_data!, along with two videos that were excluded from the final dataset but used in the challenge for reproducibility.
 The dataset is available for non-commercial use under a CC BY license and requires citation of the Appendix300 publication.
@@
 Kolbinger, from the Department of Visceral, Thoracic and Vascular Surgery, Faculty of Medicine and University Hospital Carl Gustav Carus of the TUD Dresden University of Technology and the Weldon School of Biomedical Engineering, Purdue University coordinated the challenge data collection from four hospitals in Germany.
-Oliver S.
-Lestner and Jakob N.
+Oliver L.
+Saldanha and Jakob N.
 Kather, also from the Else Kröner Fresenius Center for Digital Health, supported the data aggregation from the technical point of view and obtained the source data from these centers.
@@
 2 Avg San 0.4412 0.3394 0.2194 0.6176 0.5983 NaN $15\times 10^{-177} 1.1$$17\times 10^{-128}$ NaN
-Appendix H Credit Authorship Contribution Statement
+Appendix H Further Analysis Results with Baselines
+Table 11: Comprehensive ranking summary for Task 1, Task 2, and overall performance including contextual baselines. Lower rank indicates better performance.
+Team Task 1 Task 2 Overall Final
+EC F1 Avg EC F1 Avg Avg Rank
+Camma 5 5 5 5 2 3 4.00 3
+Elbflorenz 4 4 4 4 5 4 4.00 3
+Santhi 2 2 2 3 4 3 2.50 2
+Cen 1 1 1 1 1 1 1.00 1
+SWARM 3 3 3 2 3 2 2.50 2
+H.0.1 Detailed Bootstrap Analysis Tables
+Task 1: Generalization (F1-score)
+Tables 1214 report rank frequency, pairwise win probabilities, and Wilcoxon p-values for Task 1 F1-score.
+Rankings are stable: the centralized baseline holds rank 1 in 67.0% of bootstrap iterations and all pairwise comparisons are statistically significant ($p<0.001$).
+Table 12: Task 1 F1-score: Rank frequency distribution.
+Team Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
+Camma 0.0000 0.0000 0.0007 0.2486 0.7507
+Elbflorenz 0.0002 0.0022 0.0111 0.7385 0.2480
+Santhi 0.2961 0.4736 0.2233 0.0070 0.0000
+Cen 0.6700 0.2917 0.0377 0.0005 0.0001
+SWARM 0.0350 0.2319 0.7265 0.0065 0.0001
+Table 13: Task 1 F1-score: Pairwise win probabilities $P(\text{team}_{i}>\text{team}_{j})$.
+Team Camma Elbflorenz Santhi Cen SWARM
+Camma 0.2481 0.0000 0.0001 0.0007
+Elbflorenz 0.7509 0.0077 0.0011 0.0074
+Santhi 1.0000 0.9923 0.3024 0.7637
+Cen 0.9999 0.9989 0.6974 0.9330
+SWARM 0.9992 0.9926 0.2361 0.0654
+Table 14: Task 1 F1-score: Wilcoxon p-values. All comparisons $p<0.001$.
+Team Camma Elbflorenz Santhi Cen SWARM
+Camma 0.000 0.000 0.000 0.000
+Elbflorenz 0.000 0.000 0.000 0.000
+Santhi 0.000 0.000 0.000 0.000
+Cen 0.000 0.000 0.000 0.000
+SWARM 0.000 0.000 0.000 0.000
+Task 1: Generalization (Expected Cost)
+Tables 1517 report the equivalent analysis for the EC metric.
+Rankings are similarly stable, with Camma retaining the last rank in 100% of iterations and all pairwise differences statistically significant.
+Table 15: Task 1 EC: Rank frequency distribution.
+Team Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
+Camma 0.0000 0.0000 0.0000 0.0000 1.0000
+Elbflorenz 0.0000 0.0005 0.0035 0.9960 0.0000
+Santhi 0.2576 0.3401 0.4003 0.0020 0.0000
+Cen 0.7384 0.2480 0.0136 0.0000 0.0000
+SWARM 0.1021 0.3715 0.5254 0.0010 0.0000
+Table 16: Task 1 EC: Pairwise win probabilities.
+Team Camma Elbflorenz Santhi Cen SWARM
+Camma 0.0000 0.0000 0.0000 0.0000
+Elbflorenz 1.0000 0.0021 0.0001 0.0013
+Santhi 1.0000 0.9973 0.2048 0.5382
+Cen 1.0000 0.9999 0.7355 0.8858
+SWARM 1.0000 0.9983 0.4091 0.0703
+Table 17: Task 1 EC: Wilcoxon p-values. All highly significant ($p<0.001$).
+Team Camma Elbflorenz Santhi Cen SWARM
+Camma 0.000 0.000 0.000 0.000
+Elbflorenz 0.000 0.000 0.000 0.000
+Santhi 0.000 0.000 0.000 $2.89\times 10^{-65}$
+Cen 0.000 0.000 0.000 0.000
+SWARM 0.000 0.000 $2.89\times 10^{-65}$ 0.000
+Task 2: Adaptation (F1-score)
+Tables 1820 show markedly higher rank volatility compared to Task 1.
+No team retains a single rank in more than 66.4% of iterations, and the pairwise win probabilities between Camma, Santhi, and SL indicate substantial overlap, reflecting the instability of local adaptation across centers.
+Table 18: Task 2 F1-score: Rank frequency distribution.
+Team Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
+Camma 0.1794 0.3490 0.2499 0.1517 0.0700
+Elbflorenz 0.0078 0.0517 0.1407 0.3063 0.4935
+Santhi 0.0153 0.0992 0.2188 0.3254 0.3413
+Cen 0.6641 0.1999 0.0866 0.0367 0.0127
+SWARM 0.1335 0.3001 0.3041 0.1798 0.0825
+Table 19: Task 2 F1-score: Pairwise win probabilities.
+Team Camma Elbflorenz Santhi Cen SWARM
+Camma 0.8507 0.7786 0.2326 0.5540
+Elbflorenz 0.1493 0.4091 0.0408 0.1748
+Santhi 0.2212 0.5909 0.0648 0.2447
+Cen 0.7674 0.9592 0.9352 0.8042
+SWARM 0.4460 0.8252 0.7553 0.1958
+Table 20: Task 2 F1-score: Wilcoxon p-values.
+Team Camma Elbflorenz Santhi Cen SWARM
+Camma 0.000 0.000 0.000 $9.54\times 10^{-35}$
+Elbflorenz 0.000 0.000 0.000 0.000
+Santhi 0.000 $1.49\times 10^{-101}$ 0.000 0.000
+Cen 0.000 0.000 0.000 0.000
+SWARM $9.54\times 10^{-35}$ 0.000 0.000 0.000
+Task 2: Adaptation (Expected Cost)
+Tables 2123 report Task 2 EC bootstrap results, where rank instability is most pronounced.
+Rank frequencies are close to uniform for most teams, confirming that EC-based adaptation rankings are not reliably reproducible under resampling and should be interpreted with caution.
+Table 21: Task 2 EC: Rank frequency distribution.
+Team Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
+Camma 0.1367 0.1533 0.1737 0.2169 0.3194
+Elbflorenz 0.1262 0.1773 0.2202 0.2377 0.2386
+Santhi 0.2035 0.2317 0.2299 0.1956 0.1393
+Cen 0.3043 0.2228 0.1840 0.1628 0.1261
+SWARM 0.2323 0.2154 0.1933 0.1849 0.1741
+Table 22: Task 2 EC: Pairwise win probabilities.
+Team Camma Elbflorenz Santhi Cen SWARM
+Camma 0.4637 0.3791 0.3333 0.3872
+Elbflorenz 0.5336 0.4015 0.3571 0.4156
+Santhi 0.6166 0.5954 0.4415 0.5026
+Cen 0.6665 0.6422 0.5578 0.5477
+SWARM 0.6123 0.5839 0.4971 0.4517
+Table 23: Task 2 EC: Wilcoxon p-values. Note: Santhi vs. SWARM $p=0.746$ (not significant).
+Team Camma Elbflorenz Santhi Cen SWARM
+Camma $1.30\times 10^{-21}$ $2.72\times 10^{-189}$ 0.000 $1.21\times 10^{-171}$
+Elbflorenz $1.30\times 10^{-21}$ $3.81\times 10^{-72}$ 0.000 $4.71\times 10^{-98}$
+Santhi $2.72\times 10^{-189}$ $3.81\times 10^{-72}$ 0.000 0.746
+Cen 0.000 0.000 0.000 $2.86\times 10^{-38}$
+SWARM $1.21\times 10^{-171}$ $4.71\times 10^{-98}$ 0.746 $2.86\times 10^{-38}$
+Appendix I Credit Authorship Contribution Statement
 Max Kirchner: Conceptualization, Challenge Organization, Methodology, Software, Validation, Formal Analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review and Editing, Visualization, Project Administration Hanna Hoffmann: Conceptualization, Software, Writing - Review and Editing Alexander C.